# In-context Prompt Learning for Test-time Vision Recognition with Frozen Vision-language Model

### v1

@@ 3.1 Preliminaries: test-time prompt learning
3.2 Visual prompt input with images
The visual prompt approach enables the learning of image perturbations (such as noisy boxes or rectangles) for a frozen model. This allows the model, when prompted with these perturbations, to perform a new task. As depicted in Figure 3 (a) and (b), recent studies [2, 15] have introduced task-specific learnable parameters as visual prompts. These prompts are prepended into the input image space and are learned during the fine-tuning stage. However, these parameters are typically initialized with random noise. We contend that a general text description (e.g., a photo caption) provides richer visual context information compared to random noise.
Figure 3: The distinction between our language-aware visual prompt approach and the patched, padded, and token-based visual prompt methods.
This is because linguistic words can be perceived as interpretable lexical tokens that can be easily translated into visual content comprehensible to the vision encoder. This is especially advantageous during the test-time adaptation.
A straightforward approach to implementing a visual prompt is to combine patch image embedding with learnable parameter vectors initialized with word embeddings. However, this design not only lacks a seamless connection between the vision and language modalities but also faces the challenge of differences in dimensionality between these two types of embeddings. In this regard, we propose an efficient language-to-vision translation design that has proven to be highly effective in practice. Specifically, in addition to the word context vectors, we introduce the concept of a lightweight neural network, known as Token-Net. Token-Net is responsible for generating a conditional token vector for each input image, which is then integrated with the context visual vectors. For a visual representation of this architecture, please refer to Figure 2.
Let $\mathcal{F}_{\theta}$ denote token net parameterized by $\theta$, each visual prompt $\bm{P}_{\text{{v}}}$ is obtained by projecting text words in a prompt $\bm{P}_{\text{{t}}}$ into a set of learnable vectors via language-to-vision token net, such that $\bm{P}_{\text{{v}}}=f_{\theta}(\bm{P}_{\text{{t}}})$. Given as input image $\textbf{X}\in\mathbb{R}^{C\times H\times W}$, we divide it into $M$ pathes and produce patch tokens through a projection at vision branch of CLIP model. The visual prompt for the $i$-th input image is thus conditioned on the language description. The token net is built with a linear layer, which maps inputs of dimensions $d_{l}$ to $d_{v}$. We leave the exploration of more advanced designs to future endeavors. During training, we update the context vectors $\bm{P}_{\text{{v}}}$ together with Token-Net’s parameters $\mathbf{\theta}$. This acts as a bridge between visual and language modalities, thus reducing modality discrepancy and encouraging knowledge transfer from language to vision. Unlike CoCoOP [51], which solely conditions a text prompt on each input image without the learning of visual prompts, the explicit conditioning of our visual prompt on natural language ensure the initial visual prompts within a meaningful embedding space, thereby facilitating a smoother convergence process.
Figure 4: Illustration on the proposed visual in-context prompt learning for test-time visual recognition. Each in-context example ($x_{i},y_{i}$), test sample $x_{t}$, and its prefix text are fed into a token encoder to obtain visual, prefix, and text tokens. The text tokens are translated into a visual prompt using a token network. We optimize the visual tokens using a context-aware objective: a supervised cross-entropy term $L(x_{i},y_{i},\bm{P}_{\text{{v}}})$ involving in-context examples and an unsupervised entropy minimization $L(x_{t},\bm{P}_{\text{{v}}})$ with the test sample.
@@ 3.3 In-context learning for visual recognition
In-context learning is an emerging paradigm that involves supplying labeled examples as input to a pre-trained language model. Alongside these examples, a query example is provided for which the model must predict the label based on its understanding of the task gleaned from the provided examples. Unlike conventional few-shot learning methods, in-context learning does not necessitate the updating of all weight parameters of the pre-trained model or the learning of task-specific prompts with training data for each task. Our method, InCPLL, extends this approach into the realm of visual recognition. It does so by conditioning visual prompt learning on task-specific image-label pairs while keeping the model parameters frozen. As a result, our proposed method eliminates the need for customizing models for individual tasks and generates sample-aware prompts for the foundational model. The illustration of the proposed InCPLL is shown in Figure 4.
@@
In-context examples $\mathcal{D}$ offers contextual guidance (prompt) to the model $\mathcal{F}(\cdot;\Phi)$, enabling it to generate the optimal $y_{t}$ for $x_{t}$ without updating model parameters. As a kind of unsupervised prompt learning [50], in-context learning relies heavily on the pre-trained model and its objective may well not align with that used in pre-training. Our InCPLL append visual prompts given as $\bm{P}_{\text{{v}}}$ with the visual tokens $\{e_{cls},e_{1},e_{2},\ldots,e_{M}\}$. Hence, the visual encoder processes the following input tokens $\widetilde{X}_{p}=\{\bm{P}_{\text{{v}}},e_{cls},e_{1},e_{2},\ldots,e_{M}\}$ to produce the prompted visual feature represented as $f_{\widetilde{p}}=\mathcal{F}(X_{p};\Phi)$. Then, the instance-specific visual prompt is learned on the following objective to provide the model with the most helpful context information about on-the-fly test sample $x_{t}$.
@@
Here $\lambda$ is loss weight parameter balancing the contributions of different loss components. Each of these samples only necessitates a single pass through the visual-language model, coupled with a shared prompt. Following a single-step optimization process involving the context-aware objective, which incorporates on-the-fly test samples, we assess the online performance of the proposed InCPLL method on these test samples, utilizing the learned prompt. In the end, we utilize the average model performance on the testing data as the evaluation metric.
@@ 3.4 Cyclic Learning of Visual and Language Prompts
Visual and language (V-L) prompts are two parameter-effective prompt-tuning methods. Most methods [16, 17] attempt to simultaneously adapt vision and language branch of CLIP model with learnable prompts of both modalities, which ensures mutual synergy between V-L prompts. This mutual learning strategy performs effectively when training data provides alignment information across visual and language modalities. However, this performance is limited for the target distribution during test-time adaptation, where on-the-fly test sample is unlabeled (i.e., only visual information is available). To this end, we propose a cyclic learning of visual and language prompt to tune vision and language branch of CLIP sequentially, which leverages context information obtained from one-shot examples. Specifically, we start by optimizing one prompt and subsequently utilize the knowledge gained from this step to guide the optimization of the remaining prompt. This process provides the model with richer contextual information, allowing it to thoroughly capture relevant patterns of on-the-fly test sample to align visual and language modalities. The optimization objective for the proposed cyclic learning strategy can be formulated by
@@
$\scriptsize\mathcal{L}(x_{t},\mathcal{D},\bm{P}_{\textbf{v}},\bm{P}_{\textbf{t}},\theta)=\begin{cases}\arg\min_{\bm{P}_{\textbf{v}},\theta}\left\{L(x_{t},\bm{P}_{\textbf{v}})+\sum_{(x_{i},y_{i})\in\mathcal{D}}\lambda L(x_{i},y_{i},\bm{P}_{\textbf{v}},\bm{P}_{\textbf{t}},\theta)\right\},s=1,\\ \arg\min_{\bm{P}_{\textbf{t}}}\left\{L(x_{t},\bm{P}_{\textbf{t}})+\sum_{(x_{i},y_{i})\in\mathcal{D}}\lambda L(x_{i},y_{i},\bm{P}_{\textbf{v}},\bm{P}_{\textbf{t}},\theta)\right\},s=2,\end{cases}$ (3)
@@
where $x_{t}$ is unlabeled test sample, $\mathcal{D}$ contains in-context examples, $\theta$ is the parameter of token net, and $s=1,2$ are the step numbers of model optimization for visual and text prompt, respectively. It is noted that the paragraphs above section 3.4 are focused on visual prompt learning, which is a single-step optimization process. In this section, we introduce text prompts to integrate visual prompts. This process is a two-step optimization process, where each type of prompt is optimized in sequence, i.e., first visual then textual prompts. It forms the ”cyclic prompt learning” concept.
@@ 4 Experiments
Datasets. Here we evaluate the transfer performance of our method on 9 fine-grained classification datasets and 4 out-of-distribution (OOD) data of ImageNet. Fine-grained includes specific species of plant or animal species (Flowers102 [25], OxfordPets [26]), food (Food101 [4]), transportation (StanfordCars [18]), scenes [46], aircraft (Aircraft [22]), textures (DTD [7]), satellite images (EuroSAT [11]), human actions (UCF101 [35]), and general objects (Caltech101 [9]). OOD datasets is from ImageNet variants, i.e. ImageNet-V2 [28], ImageNet-Sketch [42], ImageNet-A [13] and ImageNet-R [12]. For more detailed information about the datasets and implementation details, we refer readers to Section 6.1 in the appendix.
@@ 4.1 Generalize CLIP to fine-grained classification datasets
In this experiment, we investigate the cross-dataset generalization ability of CLIP model from ImageNet to various fine-grained datasets. Experimental results are reported in table 2. It can be seen that all prompt learning approaches perform better than the zero-shot generalization of CLIP model. However, the learned few-shot prompts (i.e., CoOp [52], CoCoOp [51]) are constrained by the distribution corresponding to training data and may exhibit limited generalization beyond that, especially for fine-grained datasets. The learned zero-shot prompts (i.e., TPT [32]) catastrophically underfit to novel tasks, whereas our model successfully learns all tasks with the help of task-specific examples. As demonstrated in Table 2, our model significantly outperforms zero-shot and few-shot baselines.
Table 3: Comparison with existing methods across a range of distribution shifts on the ImageNet dataset. CoOp, CoCoOp, CLIPOOD, MaPLe and PromptSRC are trained on source domain (ImageNet) and evaluated on out-of-domain datasets (ImageNet variant distribution). Baseline CLIP, prompt ensemble, and TPT don’t require ImageNet dataset as training data.
@@ 4.2 Generalize CLIP to distribution Shift
In this experiment, we evaluate the out-of-distribution generalization of CLIP model on different variants of ImageNet. A particular in-distribution dataset is ImageNet, while we use several out of distribution datasets from ImageNet variants, i.e. ImageNet-V2 [28], ImageNet-Sketch [42], ImageNet-A [13] and ImageNet-R [12]. The performance on four variants of ImageNet with distribution shifts are summaried in Table 3. As shown in Table 3, CoOp [52], CoCoOp [51], CLIPOOD [33], MaPLe [16] and PromptSRC [17] are few-shot prompt tuning methods, which are trained on ImageNet using 16-shot training data per category. This is a type of rapid adaptation from ImageNet to ImageNet Variants, thus achieving better classification accuracies than zero-shot generalization of CLIP model. TPT [32] learn the language prompt using only the given test sample and enhances the zero-shot performance of CLIP model without any task-specific training data. Our InCPL extends TPT to visual prompt learning and learn task-specific information from in-context examples of testing data distribution. While our InCPL is not trained on ImageNet dataset, it inherits the zero-shot abilities of CLIP model and achieve on-par generalization as ImageNet trained baselines (CoOp [52], MaPLe [16], PromptSRC [17]), surpassing existing methods in terms of average accuracy.
Table 4: Study on task and instance-specific adaptations. w/o U/S/CU-obj: without unsupervised/supervised/context-aware unsupervised objective.
Method Adaptation Flower DTD Pets
Top 1 acc. Top 1 acc. Top 1 acc.
Ours w/o U-obj Task 69.39 42.43 87.84
Ours w/o S-obj Instance 67.24 24.05 87.11
Ours (CU-obj) Task & Instance 71.13 47.34 90.60

### v2

@@ 3.1 Visual Prompt Input with Images
Formally, let $\mathcal{F}_{\theta}$ denote token net parameterized by $\theta$, each visual prompt $\bm{P}_{\text{{v}}}$ is obtained by projecting text words in a prompt $\bm{P}_{\text{{t}}}$ into a set of learnable vectors via language-to-vision token net, such that $\bm{P}_{\text{{v}}}=f_{\theta}(\bm{P}_{\text{{t}}})$. Given as input image $\textbf{X}\in\mathbb{R}^{C\times H\times W}$, we divide it into $M$ pathes and produce patch tokens through a projection at vision branch of CLIP model. The visual prompt for the $i$-th input image is thus conditioned on the language description. The token net is built with a linear layer, which maps inputs of dimensions $d_{l}$ to $d_{v}$. We leave the exploration of more advanced designs to future endeavors. During training, we update the context vectors $\bm{P}_{\text{{v}}}$ together with Token-Net’s parameters $\mathbf{\theta}$. This works as a bridge between visual and language modalities, thus reducing modality discrepancy and encouraging knowledge transfer from language to vision. Unlike CoCoOP [49], which solely conditions a text prompt on each input image without visual prompts, the explicit conditioning of our visual prompt on natural language ensure the initial visual prompts within a meaningful embedding space, thereby facilitating a smoother convergence process.
Figure 4: Illustration of the proposed visual in-context prompt learning for test-time visual recognition. Each in-context example ($x_{i},y_{i}$), test sample $x_{t}$, and its prefix text are fed into a token encoder to obtain visual, prefix, and text tokens. The text tokens are translated into a visual prompt using a token network. We optimize the visual tokens using a context-aware objective: a supervised cross-entropy term $L(x_{i},y_{i},\bm{P}_{\text{{v}}})$ involving in-context examples and an unsupervised entropy minimization $L(x_{t},\bm{P}_{\text{{v}}})$ with the test sample.
@@ 3.2 In-context Learning for Visual Recognition
In-context learning is an emerging paradigm that involves supplying labeled examples as input to a pre-trained language model. Alongside these examples, a query example is provided for which the model must predict the label based on its understanding of the task learned from the provided examples. Unlike existing few-shot prompt learning methods, which learn share prompt for all test data, our InCPL can learn adaptive prompts tailored to different test samples. The illustration of the proposed InCPL is shown in Figure 4.
@@
Here in-context examples $\mathcal{D}$ offer contextual guidance (prompt) to the model $\mathcal{F}(\cdot;\Phi)$, enabling it to generate the optimal label $y_{t}$ for $x_{t}$ without updating model parameters. Traditional in-context learning relies heavily on the pre-trained model, while its objective on downstream task may well not align with that used in pre-training. Hence, our InCPL appends visual prompts given as $\bm{P}_{\text{{v}}}$ with the visual tokens $\{e_{cls},e_{1},e_{2},\ldots,e_{M}\}$, and tune them to learn task-specific information. Specifically, the visual encoder processes the following input tokens $\widetilde{X}_{p}=\{\bm{P}_{\text{{v}}},e_{cls},e_{1},e_{2},\ldots,e_{M}\}$ to produce the prompted visual feature represented as $f_{\widetilde{p}}=\mathcal{F}(X_{p};\Phi)$. Then, visual prompt is learned on the following objective to provide the model with the helpful context information about on-the-fly test sample $x_{t}$.
@@
Here $\lambda$ is loss weight parameter balancing the contributions of different loss components.
@@ 3.3 Cyclic Learning of Visual and Language Prompts
Visual and language (V-L) prompts are two parameter-effective prompt-tuning methods. Most methods [16, 17] attempt to simultaneously adapt vision and language branch of CLIP model with learnable prompts of both modalities, which ensures mutual synergy between V-L prompts. This mutual learning strategy performs effectively when training data provides alignment information across visual and language modalities. However, this performance is limited for the target distribution during test-time adaptation, where on-the-fly test sample is unlabeled (i.e., only visual information is available).
To this end, we propose a cyclic learning strategy for visual and textual prompts to ensure the mutual synergy across different modalities. Specifically, we start by optimizing one prompt and subsequently utilize the knowledge gained from this step to guide the optimization of the remaining prompt. This process provides the model with richer contextual information, allowing it to thoroughly capture relevant patterns of on-the-fly test sample to align visual and language modalities. The optimization objective for the proposed cyclic learning strategy can be formulated by
@@
$\scriptsize\mathcal{L}(x_{t},\mathcal{D},P_{v},P_{t},\theta)=\begin{cases}\begin{aligned} \arg\min_{\bm{P}_{\textbf{v}},\theta}\Bigg{\{}&L(x_{t},\bm{P}_{\textbf{v}})+\sum_{(x_{i},y_{i})\in\mathcal{D}}\lambda L(x_{i},y_{i},\bm{P}_{\textbf{v}},\bm{P}_{\textbf{t}},\theta)\Bigg{\}},\\ &\quad\text{if }s=1,\end{aligned}\\ \begin{aligned} \arg\min_{\bm{P}_{\textbf{t}}}\Bigg{\{}&L(x_{t},\bm{P}_{\textbf{t}})+\sum_{(x_{i},y_{i})\in\mathcal{D}}\lambda L(x_{i},y_{i},P_{v},\bm{P}_{\textbf{t}},\theta)\Bigg{\}},\\ &\quad\text{if }s=2.\end{aligned}\end{cases}$ (3)
@@
where $x_{t}$ is unlabeled test sample, $\mathcal{D}$ contains in-context examples, $\theta$ is the parameter of token net, and $s=1,2$ are the step numbers of model optimization for visual and text prompt, respectively. It is noted that the paragraphs above section 3.4 are focused on visual prompt learning, which is a single-step optimization process. In this section, we introduce text and visual prompt learning in cyclic strategy, which is is a two-step optimization process. Two kinds of prompts are optimized in sequence, i.e., first visual then textual prompts, forming the “cyclic prompt learning” concept.
3.4 Discussion
In-context learning allows large language models (GPT3 [4], Llama [39]) to perform inference on unseen tasks by conditioning on in-context examples (a.k.a. prompt) without updating the model parameters. Inspired by this, existing works [43, 48, 42] explore “in-context learning” concept for vision model, in which the model is updated using in-context examples. Meanwhile, CLIP itself is not able to conduct in-context learning task. To equip CLIP with this ability, our InCPL introduces learnable prompt for each test sample in test-time stage. In this way, the model can automatically understand the underlying task with in-context examples.
Comparison with Few-shot Learning. As few-shot methods, CoOP [50] and CoCoOP [49] fine-tune the prompt on ImageNet dataset using 16-shot training data per category and evaluate the generalization performance on downstream tasks. Our work primarily differs from few-shot methods in two main aspects, i.e., sample selection and quantity. (a) For sample selection, few-shot uses strict categories with specific number of samples, which are widely used in training stage. Differently, in-context learning has no constraint on category. The in-context samples in testing stage can either share the same category as the current test sample or the irrelevant category. It is also impractical to know the exact category of unlabeled test sample in advance. (b) For sample quantity, few-shot learning requires a predefined number of samples from each category, while in-context learning uses a small, arbitrary set of labeled samples-commonly just five samples.
Comparison with Semi-supervised Learning. Semi-supervised learning [37, 34] typically incorporates labeled data during the training phase, amalgamating it with unlabeled data to fine-tune the model and improve its performance on unlabeled samples. Labeled data in semi-supervised learning often shares categories with the unlabeled data. In our method, there is no inherent relationship between in-context examples (labeled data) and the test sample, as they are both drawn from the same domain dataset. Our approach does not necessitate any category information about the test sample, distinguishing it from semi-supervised learning methods.
@@ 4 Experiments
4.1 Datasets and Implementation details
4.1.1 Datasets
We assess the transfer performance of our method across 10 fine-grained classification datasets and 4 out-of-distribution (OOD) datasets from ImageNet. Since our method primarily focuses on test-time model adaptation, our evaluation is exclusively based on the testing dataset across all these datasets.
• Flower102 [25] is a widely used dataset consisting of 102 different categories of flowers. Each category consists of between 40 and 258 images. It is commonly employed for fine-grained image classification tasks.
• OxfordPets [26] is a dataset designed for pet image classification, containing a large variation in scale, pose, and lighting conditions. It contains images of 37 different pet breeds with roughly 200 images for each class.
• Food101 [3] is a dataset specifically curated for food recognition applications. It contains images of 101 different food categories, making it suitable for tasks related to food image classification and analysis.
• Describable Textures Dataset (DTD) [6] is a texture dataset, which consists of 5640 images. These images are classified into 47 distinct categories, inspired by human perception, with precisely 120 images allocated to each category.
• StanfordCars [18] is a dataset commonly used for fine-grained car classification tasks. This dataset contains 16,185 images of 196 classes of cars, which is split into 8,144 training images and 8,041 testing images.
• Aircraft [22] dataset contains 10,200 images of various aircraft, with 100 images for each of 102 different aircraft model variants, most of which belong to airplanes.
• UCF101 [35] is a widely recognized dataset for human action recognition, which consists of 13,320 video clips spanning 101 different human action categories. These 101 categories are further classified into 5 types, including Body motion, Human-object interactions, Playing musical instruments, Sports, and Human-human interactions.
• EuroSAT [11] is a dataset and deep learning benchmark designed for land use and land cover classification. It is based on Sentinel-2 satellite images with 13 spectral bands and a total of 27,000 labeled and geo-referenced images with 10 distinct classes.
• Caltech101 [8] dataset is composed of approximately 9,000 images with 101 object categories and a background category. Each object category contains approximately 40 to 800 images, with typical image sizes of 200-300 pixels.
• SUN397 dataset encompasses 108,753 images spanning 397 categories, serves as a benchmark in scene understanding studies. Each category in this diverse collection is represented by a minimum of 100 images.
• ImageNet [7] dataset is a large-scale ontology of images built upon the backbone of the WordNet structure, designed to advance the field of computer vision. It spans 1000 distinct object classes and contains 1,281,167 training images, 50,000 validation images and 100,000 test images.
• ImageNet-V2 [28] is a test set including natural images collected from various sources. This dataset is consisted of 10,000 images distributed across 1,000 distinct ImageNet categories.
• ImageNet-Sketch [41] dataset consists of 50000 images, 50 images for each of the 1000 ImageNet classes. These images are obtained by performing Google Image searches using the query ”sketch of [standard class name].”
• ImageNet-A [13] is a challenging dataset containing real-world, unmodified, and naturally occurring examples that are misclassified by ResNet models. It includes 7,500 intentionally altered and corrupted images with 1,000 categories to assess the robustness of different models.
• ImageNet-R [12] collects images of ImageNet categories presented in artistic renditions, including a total of 30,000 images across 200 distinct ImageNet categories.
Implementation Details. We apply in-context prompt learning on a pre-trained ViT-B/16 CLIP model. We minimize the semi-parametric objective loss to optimize the visual prompt for 1 step and cyclically alternate between optimizing the visual and prompt for 2 steps. All models are trained with in-context examples and test sample on a single NVIDIA GPU. We use the AdamW [21] optimizer with a learning rate of $5e^{-3}$ for all datasets. By default, we set weight parameter $\lambda=0.4$. If not specifically emphasized, we adopt in-context prompt learning with only visual prompt as our default setting for all ablation studies. We select 5 in-context examples from a fixed subset of labeled data, which is composed of randomly sampling 1 sample from each category. The context samples only provide the task information to do on the test sample. They are usually from the same target dataset, but there are no other relationships between them, such as category. Token net is randomly initialized at the start of test-time adaptation and accumulatively updated across the entire evaluation process. For each test sample, $P_{\text{t}}$ is initialized with prefix tokens derived from “a photo of a”, which is then converted into a visual token. $P_{\text{v}}$ is then initialized by the above learned $P_{\text{t}}$.
@@ 4.2 Generalize CLIP to Fine-grained Classification Datasets
In this experiment, we investigate the cross-dataset generalization ability of CLIP model from ImageNet to various fine-grained datasets. Experimental results are reported in table 2. It can be observed that all prompt learning approaches perform better than the zero-shot generalization of CLIP model. However, the learned few-shot prompts (i.e., CoOp [50], CoCoOp [49]) are constrained by the distribution corresponding to training data and may exhibit limited generalization beyond that, especially for fine-grained datasets. The learned zero-shot prompts (e.g., TPT [32]) catastrophically underfit to novel tasks, whereas our model successfully learns all tasks with the help of in-context information. As shown in Table 2, our model achieves the best performance over zero-shot and few-shot baselines in terms of average accuracy.
Table 3: Comparison with existing methods across a range of distribution shifts on the ImageNet dataset. CoOp, CoCoOp, CLIPOOD, MaPLe, and PromptSRC are trained on source domain (ImageNet) and evaluated on out-of-domain datasets (ImageNet variant distributions). Baseline CLIP, prompt ensemble, and TPT don’t require ImageNet dataset as training data.
@@ 4.3 Generalize CLIP to Distribution Shift
In this experiment, we evaluate the out-of-distribution generalization of CLIP model on different variants of ImageNet. A particular in-distribution dataset is ImageNet, while we use several out-of-distribution (OOD) datasets from ImageNet variants, i.e., ImageNet-V2 [28], ImageNet-Sketch [41], ImageNet-A [13], and ImageNet-R [12]. The performance on four variants of ImageNet with distribution shifts is summarized in Table 3. As shown in Table 3, CoOp [50], CoCoOp [49], CLIPOOD [33], MaPLe [16], and PromptSRC [17] are few-shot prompt tuning methods, which are trained on ImageNet using 16-shot training data per category. This is a type of rapid adaptation from ImageNet to ImageNet variants, thus achieving better classification accuracies than zero-shot generalization of CLIP model. TPT [32] learns the language prompt using only the given test sample and enhances the zero-shot performance of CLIP model without any task-specific training data. Our InCPL extends TPT to visual prompt learning and learns task-specific information from in-context examples of testing data distribution. While our InCPL is not trained on ImageNet dataset, it inherits the zero-shot abilities of CLIP model and achieves on-par generalization as ImageNet trained baselines (CoOp [50], MaPLe [16], PromptSRC [17]), surpassing existing methods in terms of average accuracy.
Table 4: Ablation studies on task and sample-specific adaptations in our InCPL. w/ U/S/CU-obj: with unsupervised/supervised/context-aware unsupervised objective.
Method Adaptation Flower DTD Pets Imagenet-R ImageNet-S
Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc.
Ours w/ S-obj Task-specific 69.36 42.43 87.84 76.89 47.24
Ours w/ U-obj Sample-specific 67.24 24.05 87.11 75.73 44.51
Ours w/ CU-obj In-context 71.13 47.34 90.60 77.56 48.03